OpenStudy (anonymous):

Solve: 10/x^2-2x x 4/x = 5/x-2

OpenStudy (anonymous):

Can you indicate where the parentheses are

OpenStudy (anonymous):

\[\frac{ 10 }{ x^{2} }\times \frac{ 4 }{ x } = \frac{ 5 }{ x-2 }\]

OpenStudy (mertsj):